### v1

@@ 1 Introduction
Molecular graph generation is pivotal for material and drug discovery, enabling efficient exploration of vast chemical spaces [Wang et al.(2025)Wang, Chen, Ma, Yu, Wang, Liu, Ye, Sakurai, and Zeng, Du et al.(2024)Du, Jamasb, Guo, Fu, Harris, Wang, Duan, Liò, Schwaller, and Blundell]. For real-world applications, a key challenge lies in controlling the generation process so that molecules satisfy multiple desired properties [Gebauer et al.(2022)Gebauer, Gastegger, Hessmann, Müller, and Schütt]. For instance, one may wish to generate a set of candidate molecules that simultaneously meet specific ranges of synthetic scores and permeability to target gases.
@@ 3.2 Multi-conditional Molecular Generation via Concrete Scores
Figure 2: An overview of CSGD with Composable Guidance (CoG) to enable flexible control ($\text{P}.$: Gas Permeability; Synth.: Synthesizability).
@@
Sample timestep $t\sim\mathcal{U}(0,1)$ and property $c_{m}\in\mathcal{C}$;
@@ 4 Experiments
To evaluate the effectiveness of extending score matching to discrete graph diffusion models, as well as the score-based capacities it enables, namely, composable guidance (CoG) and probability calibration (PC), we conducted comprehensive quantitative experiments to answer the following research questions: RQ1: Does CSGD outperform related discrete graph diffusion models? RQ2: Can concrete scores improve the quality and controllability of samples relative to the mean prediction approach? RQ3: How beneficial are CoG and PC for generating valid and controllable samples in multi-conditional generation? RQ4: Can CoG support flexible sampling under varying numbers and types of molecular property conditions?
@@ 4.3 Ablation Studies
4.5 Conclusion and Future Directions

### v2

@@ 1 Introduction
Molecular graph generation is pivotal for material and drug discovery, enabling efficient exploration of vast chemical spaces [Wang et al.(2025)Wang, Chen, Ma, Yu, Wang, Liu, Ye, Sakurai, and Zeng, Du et al.(2024)Du, Jamasb, Guo, Fu, Harris, Wang, Duan, Liò, Schwaller, and Blundell, Huang et al.(2024)Huang, Sun, Du, and Lv]. For real-world applications, a key challenge lies in controlling the generation process so that molecules satisfy multiple desired properties [Gebauer et al.(2022)Gebauer, Gastegger, Hessmann, Müller, and Schütt, Zhang et al.(2025)Zhang, Huang, Lou, Ding, Cao, and Wang]. For instance, one may wish to generate a set of candidate molecules that simultaneously meet specific ranges of synthetic scores and permeability to target gases.
@@ 3.2 Multi-conditional Molecular Generation via Concrete Scores
Condition Encoders. To handle multi-conditional generation tasks that involve both categorical and numerical properties, we adopt separate encoders for each property, following GraphDiT. Specifically, each categorical property is processed by a Categorical Encoder, and each numerical property by a Cluster Encoder. See Appendix 7.0.3 for details of these encoders. The resulting embeddings can then be combined in different ways, such as aggregating all property embeddings to form a joint representation with the mean pooling (see Figure 1C) or using a single property embedding as the input to the denoiser (see Figure 1D).
@@
Figure 2: An overview of CSGD with Composable Guidance (CoG) to enable flexible control ($\text{P}.$: Gas Permeability; Synth.: Synthesizability). 1 Dynamic Thresholding, 2 Temperature Scaling.
@@
Efficiency Concern With CoG, a separate score must be computed for each property, that is to say, the inference time increase linearly with the number of specified properties, which would be costly when this number is large. To eliminate this issue, we propose a training strategy that enables both CoG and a more efficient variant, termed fast-CoG.
During training, instead of uniformly sampling one property $c_{m}$ (Line3 in Algorithm 3.2), we randomly sample an non-empty subset $\mathcal{C}_{\text{sub}}\subseteq\mathcal{C}$ of the $M$ specified properties and aggregate their embeddings via mean pooling to form a single combined condition, which will be fed into the denoiser. The model is trained to predict the corresponding multi-conditional score $s_{\theta}(G_{t},\mathcal{C}_{\text{sub}},t)$ based on this aggregated embedding. At inference time, $s_{\theta}$, learned in this way, enables us to compute the desired multi-conditional score for any subset of specified properties with one time of model inference, which can then be directly used to make state transition. Formally, the fast-CoG score is defined as:
$s_{\theta}^{\mathrm{fast-CoG}}(G_{t},\mathcal{C}_{\text{sub}},t)=s_{\theta}(G_{t},\varnothing,t)+w\,\bigl(s_{\theta}(G_{t},\mathcal{C}_{\text{sub}},t)-s_{\theta}(G_{t},\varnothing,t)\bigr).$ (10)
Unlike CFG, fast-CoG allows flexible selection of property subsets during sampling. This training strategy enables the model to support both CoG and fast-CoG at inference, providing an option for scenarios that prioritize efficiency.
@@
Sample timestep $t\sim\mathcal{U}(0,1)$ and property $m\sim\mathcal{U}(\{1,\ldots,M\})$;
@@ 4 Experiments
To evaluate the effectiveness of extending score matching to discrete graph diffusion models, as well as the score-based capacities it enables, namely, composable guidance (CoG) and probability calibration (PC), we conducted comprehensive quantitative experiments to answer the following research questions: RQ1: Does CSGD outperform related discrete graph diffusion models? RQ2: Can concrete scores improve the quality and controllability of samples relative to the mean prediction approach? RQ3: How beneficial are CoG and PC for generating valid and controllable samples in multi-conditional generation? RQ4: Can CoG support flexible sampling under varying numbers and types of molecular property conditions? Then, we assess the sampling efficiency of CSGD with various guidance strategies to show its versatility and flexibility.
@@ 4.3 Ablation Studies
4.5 Efficiency
In the above experiments, we evaluate the performance and flexibility of CSGD using CoG. However, a natural concern arises: the inference time of CoG grows linearly with the number of specified properties, which may limit its practicality in efficiency-sensitive scenarios. To address this, in Sec. 3.2, we introduce a training strategy that enables CSGD to support both CoG and a more efficient variant, fast-CoG, simultaneously.
As shown in Table 6, we measure the average inference time per molecule on the test set using CSGD trained with that training strategy, across different property subsets and guidance strategies. The results confirm that the inference time of CoG is longer. In contrast, fast-CoG achieves comparable performance while maintaining more efficient sampling. This demonstrates the strong flexibility of CSGD: a single trained model can switch between CoG and fast-CoG with different property subsets at inference, allowing users to balance controllability and efficiency according to the needs of different scenarios.
Table 6: Evaluation of inference efficiency of CSGD using CoG and fast-CoG across different subsets of properties specified. All results are derived from the same trained model, using different sampling settings at inference. denotes the inference time per molecule, measured in seconds.
[…]
5 Conclusion and Future Directions